# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v1

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can be used as a prior for guiding the reconstruction of an undersampled subsequent contrast. To this end, several learning-based guided reconstruction methods have been proposed. However, two key challenges remain – (a) the requirement of large paired training datasets and (b) the lack of intuitive understanding of the model’s internal representation and utilization of the shared information. We propose a modular two-stage approach for guided reconstruction, addressing these challenges. A content/style model of two-contrast image data is learned in a largely unpaired manner and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased reconstruction content with clean content derived from the reference scan. We name this novel approach PnP-MUNIT. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
@@ I Introduction
Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice. However, its central limitation is the inherently slow data acquisition process. The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed. Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction. These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [5].
@@
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences and thus exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [6, 7, 8] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [9, 10, 11]. In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between the reference scan and the reconstruction.
@@
Figure 1: Our two-stage approach to guided reconstruction. (a) The first stage learns a content/style model of two-contrast MR image data in a largely unpaired manner. (b) The reconstruction stage applies this model as a content restoration operator (bottom) within an ISTA-based iterative algorithm, namely PnP-MUNIT. Guidance from the reference is introduced into the reconstruction by simply replacing its aliased content with clean content $c$ derived from the reference. CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content restoration operator. DC denotes data consistency. Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS and PnP reconstructions, respectively, are shown for comparison.
@@
The guided reconstruction problem entails using the spatial structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan. This problem has been formulated in different ways, ranging from conventional CS [6, 7] to end-to-end learning with unrolled networks [8, 12, 13] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [14]. Most learning-based approaches, although more powerful than earlier hand-crafted ones, suffer from two main limitations: (a) requirement of large paired training dataset, limiting the effective usage of real-world clinical MR data, and (b) lack of insight into the representation and utilization of the shared underlying structure. We address both issues, leveraging ideas from image-to-image modeling, content/style decomposition, and plug-and-play reconstruction. To the best of our knowledge, ours is the first work that deeply explores content/style modeling for multi-contrast MRI reconstruction.
@@
In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [15, 16, 17]. Viewed in the light of MR physics, cross-contrast translation takes an extreme stance by completely ignoring fundamental differences in contrast-specific sensitivities and relying solely on the prior contrast. Cross-contrast translation, by itself, is not a practically reliable solution because data acquisition is necessary for obtaining new information about the anatomy. That said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [18, 19], which can be adopted to complement image reconstruction. We observe that methods such as MUNIT [19] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for spatially aligned paired training data.
@@
Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging. The main research line [20, 21] has focused on learning CNN-based denoising models and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery. An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc. With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
We propose a two-stage approach for guided reconstruction. We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data. This learned model is subsequently embedded into an iterative reconstruction framework resulting in a content/style-based PnP-like algorithm [22]. An overview of our approach is shown in Fig. 1. Specifically, our contributions are four-fold:
@@
    We show that unpaired training can be used to disentangle contrast-independent and contrast-specific information as content and style, respectively, followed by a fine-tuning strategy to refine the content representation using a modest amount of paired data.
@@
    With this content/style model as the basis, we define a content restoration operator capable of removing severe undersampling artifacts from the reconstruction image given the corresponding reference image.
@@
    Developing this idea further, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
@@
    Through comprehensive experimentation, we shed light on several properties of PnP-MUNIT such as interpretability and convergence as well as demonstrating its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
@@ II-A Reconstruction Methods for Accelerated MRI
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [6] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on edge features in the reference scan. On the other hand, unrolled network-based guided methods [8, 12, 27, 28] supply the reference scan as an additional input to the reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proved to be more effective than conventional algorithms. However, they suffer from two issues limiting their practicality. First, they require large paired training datasets, which is too strong a constraint when working with retrospectively collected clinical data which reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Second, most of these methods neither explicitly represent the prior information nor provide a clear mechanism by which it is incorporated into the reconstruction. On the other hand, a conventional CS approach like STV precisely defines the prior information and the guidance mechanism. A reconstruction method combining this form of transparency with the effectiveness of deep learning-based methods, while maintaining lenient data requirement, is still needed.
@@ II-B Unpaired Image-to-Image Modeling
Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [29] and CycleGAN [30] in paired and unpaired settings, respectively. Another line of unpaired image translation methods, the first of which was UNIT [18], assumes a shared latent space underlying the two domains to explicitly represent shared information, thereby making it more interpretable than CycleGAN which directly models the image-to-image mapping. However, both assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other. Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [15, 31, 16]. However, they fundamentally do not account for the variability of the scanning setup which influence the realized contrast level and the differential visibility of pathologies in the target image. Denck et al. [17] partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level. However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e. global contrast level) in the data, for which the labels (i.e. sequence parameters) must be available.
MUNIT [19] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, overcoming the deterministic mapping limitation of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. We argue that MUNIT, due to the minimal assumptions it makes, can be generally applied to learn domain-independent anatomical structure and domain-specific variability from an unpaired multi-contrast dataset as content and style, respectively.
@@ II-C Combining Image Translation with Reconstruction
Acknowledging the fundamental limitations of MR contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of guided reconstruction involves generating a synthetic image from the reference scan via deterministic image translation and using it in L2-regularized classical reconstruction [32]. More recent work by Xuan et al. [33] proposes a joint image translation and reconstruction method based on optimal transport which additionally accounts for misalignment between the reference and reconstruction images. Levac et al. [14] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, an alternative general approach leveraging semantic content/style modeling in multi-contrast reconstruction could potentially provide a connection with the well-established MR physics concepts of quantitative tissue maps, acquisition parameters, and physical models. Such an interpretable and unifying approach remains to be found.
@@ III-A2 Plug-and-play denoiser
III-A3 Plug-and-play content restoration operator
In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the reconstruction image. Inspired by the PnP design, we cast the problem of incorporating prior information from $x^{\mathrm{ref}}_{1}$ into reconstruction $x_{2}^{k-1}$ as the restoration of its semantic content, which has been corrupted by undersampling. We propose a content restoration operator $g_{M}(.;c)$ such that (2) becomes
@@ III-A3 Plug-and-play content restoration operator
where a content/style model $M$ is used to decompose $x_{2}^{k-1}$ into content and style, followed by a replacement of this content with clean content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the image $r_{2}^{k}$ from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.

### v2

@@ Abstract
Since multiple MRI contrasts of the same anatomy contain redundant information, one contrast can be used as a prior for guiding the reconstruction of an undersampled subsequent contrast. To this end, several learning-based guided reconstruction methods have been proposed. However, a key challenge is the requirement of large paired training datasets comprising raw data and aligned reference images. We propose a modular two-stage approach for guided reconstruction addressing this issue, which additionally provides an explanatory framework for the multi-contrast problem in terms of the shared and non-shared generative factors underlying two given contrasts. A content/style model of two-contrast image data is learned from a largely unpaired image-domain dataset and is subsequently applied as a plug-and-play operator in iterative reconstruction. The disentanglement of content and style allows explicit representation of contrast-independent and contrast-specific factors. Based on this, incorporating prior information into the reconstruction reduces to simply replacing the aliased content of the image estimate with high-quality content derived from the reference scan. Combining this component with a data consistency step and introducing a general corrective process for the content yields an iterative scheme. We name this novel approach PnP-MUNIT. Various aspects like interpretability and convergence are explored via simulations. Furthermore, its practicality is demonstrated on the NYU fastMRI DICOM dataset and two in-house multi-coil raw datasets, obtaining up to 32.6% more acceleration over learning-based non-guided reconstruction for a given SSIM. In a radiological task, PnP-MUNIT allowed 33.3% more acceleration over clinical reconstruction at diagnostic quality.
@@ I Introduction
Magnetic resonance imaging (MRI) is an invaluable medical imaging modality due to the high-quality scans it delivers, the variety of complementary information it can capture, and its lack of radiation-related risks, leading to its wide usage in clinical practice. However, its central limitation is the inherently slow data acquisition process. The raw sensor data is acquired in the frequency domain, or k-space, from which the image is reconstructed. Over the last 25 years, advancements such as parallel imaging [1, 2], compressed sensing (CS) [3], and deep learning reconstruction [4, 5] have enabled considerable speedups by allowing sub-Nyquist k-space sampling and relying on computationally sophisticated reconstruction. These techniques have subsequently been implemented on commercial MRI systems and have demonstrated (potential) improvements in clinical workflow [6].
@@
A clinical MRI session typically involves the acquisition of multiple scans of the same anatomy through the application of different MR pulse sequences, exhibiting different contrasts. Because these scans are different reflections of the same underlying reality, they share a high degree of shared structure. However, currently deployed clinical protocols acquire and reconstruct each scan as an independent measurement, not leveraging the information redundancy across scans. There is, therefore, an opportunity to further optimize MRI sessions by exploiting this shared information. On the reconstruction side, multi-contrast methods have addressed this problem by introducing the shared information into the reconstruction phase to allow higher levels of k-space undersampling. In the simplest case of two contrasts, multi-contrast reconstruction can be classified into two types – (a) guided reconstruction, where an existing high-quality reference scan is used to guide the reconstruction of an undersampled second scan [7, 8, 9] and (b) joint reconstruction, where both contrasts are undersampled and are reconstructed simultaneously [10, 11, 12]. In this work, we consider the problem of guided reconstruction, assuming no inter-scan motion between reference and target scans.
@@
Figure 1: Our two-stage approach to guided reconstruction. (a) The first stage learns a content/style model of two-contrast MR image data. The two image domains $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ are decomposed into a shared content domain $\mathcal{C}$ and separate style domains $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. This illustration only provides an intuition about the information encoded in the content and style domains, and does not represent the actual training process, which is given in detail in Section III-B. (b) The reconstruction stage applies this model as a content consistency operator (bottom) within an ISTA-based iterative algorithm. Given an aligned reference image, guidance is introduced into the reconstruction by simply replacing its aliased content with content derived from the reference. CR denotes a content refinement update, which iteratively corrects for inconsistencies between the reference content and the measured k-space data, improving the effectiveness of the content consistency operator. DC denotes data consistency. Wavelet-domain soft-thresholding (top) and CNN-based denoising (middle) used in CS-WT and PnP-CNN reconstruction, respectively, are shown for comparison.
@@
The guided reconstruction problem entails using the local structure of the reference scan as a prior to complement the undersampled k-space measurements of the target scan. This problem has been formulated in different ways, ranging from conventional CS [7, 8] to end-to-end learning with unrolled networks [9, 13, 14, 15] and, more recently, diffusion model-based Bayesian maximum a posteriori estimation [16]. Most end-to-end approaches, although more powerful than earlier hand-crafted ones, suffer from the main drawback of requiring large paired training datasets consisting of the target image and its k-space together with an aligned reference image, thereby limiting their application on real-world MR data. We address this issue by proposing a plug-and-play reconstruction method that splits the problem into a purely image-domain learning sub-problem and an iterative reconstruction sub-problem. The learning problem leverages ideas from content/style decomposition, thereby offering a degree of interpretability to our approach. To the best of our knowledge, ours is the first work that thoroughly explores content/style-based generative modeling for multi-contrast MRI reconstruction.
@@
In recent years, image-to-image translation has found application in the direct estimation of one MR contrast from another [17, 18, 19, 20]. While these methods are attractive due to their lack of dependence on k-space data, viewed in the light of MR physics, cross-contrast translation takes an extreme stance by not explicitly taking into account contrast-specific sensitivities and relying solely on the prior contrast. Hence, in the context of MR image formation, cross-contrast translation, by itself, can only provide a part of the information about the target image as data acquisition via MR sequences is necessary to obtain new information about the anatomy. That being said, literature on unpaired image translation provides a repertoire of useful tools such as joint generative modeling of two-domain image data [21, 22], which can be adopted to complement image reconstruction. We observe that methods such as MUNIT [22] can be applied to learn semantically meaningful representations of contrast-independent and contrast-specific information as content and style, respectively, without the need for paired image-domain training data.
@@
Plug-and-play (PnP) methods are an emerging paradigm for solving inverse problems in computational imaging. The main research line [23, 24] has focused on learning CNN-based denoising models on image-domain data and applying them as functions replacing proximal operators in iterative algorithms like ISTA and ADMM, demonstrating improved image recovery. An advantage of this approach is the decoupling of the learning problem of image modeling from the inverse problem of image reconstruction, thereby simplifying model training and improving generalizability across different acceleration factors, undersampling patterns, etc. With this design pattern in mind, we combine content/style image modeling and iterative reconstruction in a PnP-like framework.
We first leverage semantic content/style modeling to learn explicit representations of contrast-independent and contrast-specific components from two-contrast MR image data. This training process is independent of the reconstruction problem and can be performed using unpaired images. We then make the interesting observation that in multi-contrast MR images, the style information tends to localize in the center of the k-space, allowing for an accurate style estimation from an undersampled image. Since the content of the reconstructing image, which is the remaining piece of information and is contrast-independent, is supplied by the reference image, one can compose a de-aliased estimate of the reconstruction from the undersampled image in a single step. We term this the content consistency operation, which forms the basis of our iterative reconstruction algorithm PnP-MUNIT [25]. An overview of our approach is shown in Fig. 1. While the PnP-based decoupling of image modeling and image reconstruction simplifies the training process and allows the two stages to be analyzed separately, a further level of decoupling offered by content/style disentanglement offers additional modularity and an intuitive guidance mechanism for the reconstruction.
Specifically, our contributions are four-fold:
@@
    We show that unpaired image-domain training can be used to learn disentangled contrast-independent and contrast-specific representations, followed by a fine-tuning strategy that refines the content representation using a modest amount of paired image-domain data.
@@
    With this content/style model as the basis, we define a content consistency operator capable of removing severe undersampling artifacts from the reconstructing image, given the corresponding reference image.
@@
    Developing this idea further and incorporating a corrective process for the content, we propose PnP-MUNIT, a modular algorithm for guided reconstruction, combining the flexibility of the plug-and-play approach with the semantic interpretability of content/style decomposition.
@@
    Through comprehensive experiments, we shed light on several properties of PnP-MUNIT such as convergence and robustness and demonstrate its applicability on real-world raw data and its potential clinical utility for a specific radiological task.
@@ II-A Reconstruction Methods for Accelerated MRI
One of the earliest CS-based guided reconstruction methods was proposed by Ehrhardt and Betcke [7] introducing structure-guided total variation (STV), which assumes the sparse coefficients of the reconstruction to be partially known based on the edge features in the reference scan. Later work introduced adaptive elements into the multi-contrast CS framework, e.g. Weizman et al. [8] proposed adaptive weighting-based guided CS and Song et al. [29] used adaptive sparse domains based on coupled dictionary learning. In the latter, the problem was formulated using a patch-level linear model comprising coupled and distinct sparse dictionary representations of the two contrasts. This model resembles a content/style model in form, although it is more restrictive. A general drawback of classical methods compared to deep learning-based ones is their lower flexibility. End-to-end learning-based methods [9, 13, 30, 31, 32, 15], on the other hand, supply the reference scan as an additional input to a deep reconstruction model allowing it to automatically learn the suitable features to extract and transfer into the reconstruction, and have proven to be more effective than conventional algorithms. However, end-to-end methods require large paired training datasets comprising the ground-truth image and the k-space together with aligned reference images, which is too strong a constraint when working with retrospectively collected clinical data that reflects the natural inconsistencies of routine practice such as missing contrasts, differences in spatial resolutions, and the existence of inter-scan motion. Additionally, by relying on end-to-end-learned features, these methods are generally less interpretable than their conventional counterparts in that they do not explicitly model the multi-contrast problem in terms of the underlying shared and non-shared information. MC-VarNet [15] is an important exception, which uses a simple linear decomposition of the reference contrast into common and unique components, applying the common component for guidance.
A reconstruction method that is as effective as the learning-based methods while having more lenient data requirements and offering a high degree of interpretability is still needed. Our plug-and-play method relies on learning a deep non-linear content/style transform from image-domain data only, even if subject-wise paired images are not fully available.
@@ II-B Unpaired Image-to-Image Modeling
Image-to-image modeling is the general problem of learning a mapping between two image domains and was first addressed by Pix2Pix [33] and CycleGAN [34] in paired and unpaired settings, respectively. Another line of unpaired image translation methods, the first of which was UNIT [21], assumes a shared latent space underlying the two domains to explicitly represent shared information. However, both CycleGAN and UNIT assume a deterministic one-to-one mapping between the two domains, ignoring the fact that an image in one domain can have multiple valid renderings in the other. Deterministic image translation has been widely applied to MRI contrast-to-contrast synthesis [17, 35, 18]. However, fundamentally, these methods do not account for the variability of the scanning setup that influences the realized contrast level and the differential visibility of pathologies in the target image. Denck et al. [19] partially address this problem by proposing contrast-aware MR image translation where the acquisition sequence parameters are fed into the network to control the output’s contrast level. However, this model is too restrictive since it assumes a single pre-defined mode of variability (i.e. global contrast level) in the data, for which the labels (i.e. sequence parameters) must be available.
MUNIT [22] extended UNIT by modeling domain-specific variability in addition to domain-independent structure, enabling many-to-many mapping and thus overcoming the rigidity of UNIT. The result was a stochastic image translation model which, given an input image, generates a distribution of synthetic images sharing the same “content” but differing in “style”. More fundamentally, MUNIT is a learned invertible transformation between the image domains and the disentangled content/style domains. And unlike other content/style modeling frameworks such as [36, 37, 38], MUNIT models content and style as latent generative factors of the two image domains, providing precise distribution-level definitions for them. Clinical MR images of a given protocol contain multiple modes of variability, many of which are not known a priori. Compared to [19], we make a broader assumption that the contrast-independent semantic information is local in nature and that the contrast-specific variations in the dataset can include global effects of acquisition settings as well as local anatomical features unique to the contrast. We model these using MUNIT, referring to the shared and non-shared components as content and style, respectively.
@@ II-C Combining Image Translation with Reconstruction
Acknowledging the limitation of MR cross-contrast prediction, some prior work has attempted combining it with multi-contrast reconstruction. A naive form of joint image synthesis and reconstruction, e.g. PROSIT [39] and rsGAN [40], involves generating a synthetic image from the reference scan via deterministic image translation and using it in a classical L2-regularized least-squares reconstruction. More recent work by Xuan et al. [41] proposed a joint image translation and reconstruction method which additionally accounts for misalignment between the reference and reconstructing images, with a follow-up work leveraging optimal transport theory [42]. Levac et al. [16] formulate guided reconstruction as a highly general Bayesian maximum a posteriori estimation problem and solve it iteratively via Langevin update steps, using an image-domain diffusion model as the score function of the prior distribution. While these are promising directions, we propose an alternative approach that decomposes the multi-contrast problem in a more intuitive way – first, into two sub-problems, namely image modeling and image reconstruction; and second, within the image-domain model, the multi-contrast information is decomposed into content and style. This two-level decomposition results in a highly modular reconstruction algorithm with a built-in explanatory framework where (a) the guidance mechanism is a simple content-replacement operation, (b) the discrepancy between the supplied reference content and the true content of the target image represents a meaningful error term which can be minimized, and (c) the optimal content-encoding capacity of the model for a given two-contrast image dataset indicates the amount of shared structure available to be learned in this data and used in the reconstruction task.
@@ III-A2 Plug-and-play denoiser
III-A3 Plug-and-play content consistency operator
In guided reconstruction, a spatially aligned reference $x^{\mathrm{ref}}_{1}$ is available, which captures the same underlying semantic content as the target reconstruction. Inspired by the PnP design, we cast the problem of incorporating prior information from $x^{\mathrm{ref}}_{1}$ into the reconstruction iterate $x_{2}^{k-1}$ as enforcing a hard consistency between this image iterate and its semantic content estimated from the reference. We propose a content consistency operator $g_{M}(\cdot;c)$ such that
@@ III-A3 Plug-and-play content consistency operator
where a content/style model $M$ decomposes $x_{2}^{k-1}$ into content and style, followed by a replacement of this corrupted content with high-quality content $c$ derived from $x^{\mathrm{ref}}_{1}$ and composing the improved image $r_{2}^{k}$ from it. Before formally defining this operator and developing the reconstruction algorithm, we discuss the steps required to learn the content/style model $M$.